Module 9 q3Consider the following pairwise comparison matrix for the price of three car models. Price Accord Taurus Cruz 1 Accord 1 1 3Please put answer in BOLD. Will Upvote.
Answer #1

Given table : Accord Taurus Accord 1/5 Taurus 5 Cruz 3 cruz 1/3 2 By converting above table values in fractions into normal vHence we got cruz as top priority with value of 0.500. Taurus is and priority with value of 0.376 Accord is 3rd priority withNow divide above weighted matrix by priority value we get Accord = (0.613 +0.116) = 5.28 Taurus = (2.934 + 0.376) = 7.80 Cruz